OpenStudy (anonymous):

Alexandra is two years older than brad. Cecilia is three years younger than Alexandra. All three ages added together gives a total of 37 years. How old is each student?

OpenStudy (ajspeller):

A = Alexandra B = Brad C = Cecilia A = 2 + B C = A - 3 A + B + C = 37 Solve this system of equations.